OpenStudy (wyattp17):

Calculate the length of AC given AB = 2, AE = 13.3, and AD = 1.5. Round to the nearest whole number.

OpenStudy (wyattp17):



OpenStudy (chris911):

AB = 2CM AE= 13.3 CM AD = 1.5 CM DE = AE - AD = 13.3 - 1.5 = 11.8 CM DE = BC therefore , we have AC = AB + BC = 2 + 11.8 CM = 13.8 CM so , the length of AC = 13.8 cm

OpenStudy (chris911):

do u have the multiple chocice answers?

OpenStudy (wyattp17):

no